# Characterization of a radiation detector based on opaque water-based liquid scintillator

### v1

@@ 3.3 Gamma ray response
To assess the response of the detector to ionizing radiation, gamma-ray sources were used. In all cases, the gamma-ray sources were low activity ($<5\leavevmode\nobreak\ \mu$Ci), and were placed directly on top of the lid of the detector body, centered in the horizontal plane. To a good approximation, the energy deposited in the detector by electrons is linearly proportional to the amount of emitted light. In scintillators, it is generally desirable for the signal measured to be proportional to the amount of light created. However, the amount of light collected in a channel also strongly depends on the event proximity to the fiber corresponding to that channel. The summed signal from all channels carries the information on total energy deposited; however, that information is convolved with the position-dependent light collection efficiency. Figure 8 shows the integral detector response for gamma rays from three radioisotopes, for each of the formulations of oWbLS. The general trend is that increased scattering in the liquid increases light collection as it increases the probability of the photon reaching the fiber. The shapes of the spectra also differ between the various scintillation liquids, due to their different optical properties. oWbLS1, the most transparent of the formulations, has well-defined Compton features, whereas the more highly scattering formulations (like oWbLS2), have Compton features that are broadened.
Figure 8: Response to gamma rays from three radioisotopes for (a) oWbLS1, (b) oWbLS2, and (c) oWbLS3. Among these, oWbLS1 had the longest reduced scattering length, leading to the sharpest Compton features, while oWbLS2, with the shortest reduced scattering length, provided optimal light collection. (d) Response of the three formulations of oWbLS to gamma rays from 60Co for ease of comparison. It is evident that the light collection efficiency increases as the reduced scattering length decreases.
@@ 3.5 Topological reconstruction
We investigated the use of channel-by-channel signals for an event to reconstruct event topology. The reconstructions were based on the data from the laser-pulse injection experiments, in order to have “ground-truth” values for vertex position and event intensity. The simplest method to reconstruct the event position is to calculate a quantity analogous to the center of mass (CoM) for all channels. For instance, in the $z$ direction:
@@
This experiment was repeated for four positions in the detector with two pulse intensities. The large pulses had an intensity of $20000\pm 2000$ photons/pulse, which correlates to approximately 1.6 MeV of energy deposition in oWbLS. The small pulses had an intensity of $10000\pm 1000$ photons/pulse, which is equivalent to 0.8 MeV of energy deposited in the detector. The four positions, along with their nominal offsets from the center of the detector, are listed in Table 3. The results of CoM analysis are shown in Fig. 13. The position reconstruction error is independent of the laser injection position within the fiducial region, indicating a homogeneous performance.
@@ 4 $20\pm 5$ $-20\pm 5$ $0\pm 5$ $28\pm 9$
Figure 13: Position reconstruction errors for the detector positions listed in Tab. 3. The large pulses had an intensity of $20000\pm 2000$ photons/pulse, and the small pulses were $10000\pm 1000$ photons/pulse. The average reconstruction error (shown as dotted horizontal lines) for the large pulses was 4.4 mm and for the small pulses was 7.4 mm. The horizontal error bars represent the uncertainty in offset from the center of the detector, and the vertical error bars show the standard deviation of the reconstruction position errors.
@@ 3.6 Energy Resolution
This measurement does not account for the intrinsic energy resolution of the scintillation liquid, which arises from the statistical fluctuation in the number of scintillation photons produced at a given energy deposition. The intrinsic energy resolutions for various scintillation liquids based on LAB and pseudocumene were found to vary in the range of 1.3–3.3% [Smirnov2023NoteScintillators]. While a similar measurement has not been done for oWbLS, the contribution of oWbLS to the obtained energy resolution is likely in the same range. Using laser pulses to determine energy resolution also introduces broadening as a result in the variance in intensity of the laser pulses The laser used in this experiment has a pulse-to-pulse stability of $<5$ [BorisKhouryNKTPhotonics2024PrivateCorrespondence], which partially offsets the error related to the intrinsic energy resolution of oWbLS.
@@ 4.1 Discussion
Segmenting the fiducial region of the cube with voxels of this size would require approximately 265 voxels. The 7.4 mm position reconstruction error obtained with events corresponding to 0.8 MeV could be achieved with approximately 56 voxels, or about double the number of channels used in this prototype. Moreover, this technique allows an enormous advantage when scaling up the detector volume. For example, adding another ring of fibers, (i.e., a $6\times 6$ grid as opposed to the current $4\times 4$ per side) to create a fiducial volume of approximately 1 L would require an additional 40 channels, but approximately another 1000 physically segmented voxels of side length 9.17 mm.
3D reconstruction can also be achieved with a detector that employs 2D voxels (rods). The position reconstruction in the third dimension may be accomplished by comparing the time of flight for the scintillation photons [Sutanto2021SANDD:Scintillator] at each end of the rods. Since the readout array is laid out in two planes, the number of channels scales similarly with volume to the detector described in this study. Furthermore, the energy resolution of such a system can reach 18-22% at the 137Cs Compton edge, which is superior to the results shown in Sec. 3.6. However, there are two disadvantages of the rod technique compared to an opaque scintillator. First, sub-voxel position resolution is generally not achievable in the dimensions orthogonal to the rods, whereas sub-fiber pitch resolution has been demonstrated here for opaque scintillators. And secondly, fast timing electronics are necessary to reconstruct the position along the rods. While timing information may be useful in other opaque scintillator designs, it was not used here beyond the correlation of pulses into events.
@@ 4.2 Future work
There are many opportunities to extend this work. First, an apparatus to optically characterize opaque liquid scintillators should be developed, potentially by imaging the diffuse reflectance pattern on the surface of the liquid in response to an incident pencil beam of light, as is done in Refs. [Kienle1997ImprovedMedium, Juttula2014InstrumentAngle]. Machine learning for event topological reconstruction could be appropriate for this application. One such method is to train a neural network by simulated data to generate maximum likelihood estimations of event characteristics given a series of observables [Eller2023APrinciples] (in this case, the number of photoelectrons in each channel). The detector presented above was not optimized for a particular application by adjusting the scintillator composition and doping. The quality of the topological reconstruction could almost certainly be improved by tuning the positioning and pitch of the WLS fibers. Using SiPMs instead of a single MAPMT could have at least two benefits, including a higher quantum efficiency and the ability to place the photosensors right on the outside of the detector volume, which would reduce the length and bending of the WLS fibers, which would, in turn, reduce light loss. Many materials can be investigated for opaque scintillation applications. For example, plastics and pressed organic polycrystals can be made to be highly scattering [Polupan2023PeculiaritiesPropagation]. Additionally, the reduction in the transparency requirements for detectors that use this technology raises exciting possibilities for detector doping at unprecedented levels with, for example, nanoparticles that include material with a high atomic number to increase both light scattering and the photoelectric absorption of gamma rays [Taylor2011NanofluidCollectors, Mutreja2023HighAspects]. Lastly, the base solvent used in oWbLS (DIN) has been shown to have the ability to discriminate between neutron and gamma-ray scatters based on the difference in the resulting pulse shape. The feasibility of this approach in the opaque regime should be investigated.

### v2

@@ 3.3 Gamma ray response
To assess the response of the detector to ionizing radiation, gamma-ray sources were used. In all cases, the gamma-ray sources were low activity ($<5\leavevmode\nobreak\ \mu$Ci), and were placed directly on top of the lid of the detector body, centered in the horizontal plane. To a good approximation, the energy deposited in the detector by electrons is linearly proportional to the amount of emitted light. In scintillators, it is generally desirable for the signal measured to be proportional to the amount of light created. However, the amount of light collected in a channel also strongly depends on the event proximity to the fiber corresponding to that channel. The summed signal from all channels carries the information on the total energy deposited; however, that information is convolved with the position-dependent light collection efficiency. Figure 9 shows the integral detector response for gamma rays from three radioisotopes, for each of the formulations of oWbLS. The general trend is that increased scattering in the liquid improves light collection as it increases the probability of a photon reaching the fiber. The spectral shapes also differ among various scintillators due to their different optical properties. oWbLS1, the most transparent of the formulations, has well-defined Compton features, whereas the more highly scattering formulations (like oWbLS2), have Compton features that are broadened. This is because the dependence of collection efficiency on event distance to a fiber increases as the reduced scattering length decreases.
Figure 9: Response to gamma rays from three radioisotopes for (a) oWbLS1, (b) oWbLS2, and (c) oWbLS3. Among these, oWbLS1 had the longest reduced scattering length, leading to the sharpest Compton features, while oWbLS2, with the shortest reduced scattering length, provided optimal light collection. (d) Response of the three formulations of oWbLS to gamma rays from 60Co for ease of comparison. The light collection efficiency increases as the reduced scattering length decreases. The 511 keV gamma ray from 22Na has the lowest energy of any from the radioisotopes shown. The maximum energy deposited by a single Compton scatter of a 511 keV gamma ray is approximately 340 keV, so the detector is sensitive to energy depositions of events with energy at least this low.
@@ 3.5 Topological reconstruction
We investigated the use of channel-by-channel signals for an event to reconstruct event topology. The reconstructions were based on the data from the laser-pulse injection experiments to obtain “ground-truth” values for vertex position and event intensity. The simplest method to reconstruct the event position is to calculate a quantity analogous to the center of mass (CoM) for all channels. For instance, in the $z$ direction:
@@
This experiment was repeated for four positions in the detector with two pulse intensities. The large pulses had an intensity of $20000\pm 2000$ photons/pulse, which correlates to approximately 1.6 MeV of energy deposition in oWbLS. The small pulses had an intensity of $10000\pm 1000$ photons/pulse, which is equivalent to 0.8 MeV of energy deposited in the detector. The four positions, along with their nominal offsets from the center of the detector, are listed in Table 3. The results of CoM analysis are shown in Fig. 14. The position reconstruction error is independent of the laser injection position within the fiducial region, indicating a homogeneous performance. The reconstruction error of the $y$-direction was higher than the error of the $x$-direction for the measurement taken at $(0,-8,0)$ mm, indicating that there may be small inhomogeneities within individual cells in the fiducial region. This effect appears to be minor, as the overall reconstruction error at this position was similar to the others, but should be studied in more detail in future prototypes.
@@ 4 $20\pm 5$ $-20\pm 5$ $0\pm 5$ $28\pm 9$
Figure 14: Position reconstruction errors for the detector positions listed in Table 3. The large pulses had an intensity of $20000\pm 2000$ photons/pulse, and the small pulses were $10000\pm 1000$ photons/pulse. The average reconstruction error (shown as dotted horizontal lines) for the large pulses was 4.4 mm and for the small pulses was 7.4 mm. The horizontal error bars represent the uncertainty in offset from the center of the detector, and the vertical error bars show the standard deviation of the reconstruction position errors.
@@ 3.6 Energy Resolution
This measurement does not account for the intrinsic energy resolution of the scintillation liquid, which arises from the statistical fluctuation in the number of scintillation photons produced at a given energy deposition by a charged particle. The intrinsic energy resolutions for various scintillation liquids based on LAB and pseudocumene were found to vary in the range of 1.3–3.3% [Smirnov2023NoteScintillators]. While a similar measurement has not been done for oWbLS, the contribution of oWbLS to the obtained energy resolution is likely in the same range. Using laser pulses to determine energy resolution also introduces broadening as a result in the variance in intensity of the laser pulses The laser used in this experiment has a pulse-to-pulse stability of $<5$ [BorisKhouryNKTPhotonics2024PrivateCorrespondence], which partially offsets the error related to the intrinsic energy resolution of oWbLS.
@@ 4.1 Discussion
Segmenting the fiducial region of the cube with voxels of this size would require approximately 265 voxels. The 7.4 mm position reconstruction error obtained with events corresponding to 0.8 MeV could be achieved with approximately 56 voxels, or about double the number of channels used in this prototype. It should be explicitly noted that the position resolution achieved by a physically segmented detector does not depend upon the number of photons produced by an event (above some minimum threshold), whereas it does in an opaque scintillator. Moreover, this technique allows an enormous advantage when scaling up the detector volume. For example, adding another ring of fibers, (i.e., a $6\times 6$ grid as opposed to the current $4\times 4$ per side) to create a fiducial volume of approximately 1 L would require an additional 40 channels, but approximately another 1000 physically segmented voxels of side length 9.17 mm. This analysis assumes a single, point-like interaction in the detector. Physically segmented detectors may be superior for distinguishing, for instance, two simultaneous events that take place near each other, but with sufficient separation to be located in two voxels. Additional studies are needed to determine the minimum separation between events that can be resolved in an opaque scintillation detector.
3D reconstruction has also been demonstrated with a detector that employs 2D voxels (rods). The rods are rectangular columns of transparent plastic scintillator that are optically separated from each other. Position reconstruction in the third dimension may be accomplished by comparing the amount of light collected or the relative time of arrival of scintillation photons [Sutanto2021SANDD:Scintillator] at the two ends of a rod. Since the readout array is laid out in two planes, the number of channels scales similarly with volume to the detector described in this study. Furthermore, the energy resolution of such a system can reach 18-22% at the 137Cs Compton edge, which is superior to the results shown in Sec. 3.6. However, there are two disadvantages of the rod technique compared to an opaque scintillator. First, sub-voxel position resolution is generally not achievable in the dimensions orthogonal to the rods, whereas sub-fiber pitch resolution has been demonstrated here for opaque scintillators. And secondly, fast timing electronics are necessary to reconstruct the position along the rods. While timing information may be useful in other opaque scintillator designs, it was not used here beyond the correlation of pulses into events.
@@ 4.2 Future work
There are many opportunities to extend this work, which can be grouped into three categories.
The first category relates to the optimization of detector design. This prototype detector was not optimized for a particular application by adjusting the scintillator composition and doping. The quality of topological reconstruction could almost certainly be improved by tuning the positioning and pitch of WLS fibers. Additionally, using SiPMs instead of a single MAPMT could yield at least two benefits: a higher quantum efficiency and the ability to place photosensors close to the edge of detector volume, which would reduce the length and bending of the WLS fibers and, in turn, reduce light loss.
The second category relates to opaque scintillator media. An apparatus to optically characterize opaque liquid scintillators should be developed, potentially by imaging the diffuse reflectance pattern on the surface of the liquid in response to an incident pencil beam of light, as is done in Refs. [Kienle1997ImprovedMedium, Juttula2014InstrumentAngle]. Many media can be investigated for opaque scintillation detector development. For example, plastics and pressed organic polycrystals can be made to be highly scattering [Polupan2023PeculiaritiesPropagation]. Additionally, the reduction in transparency requirements for detectors that use this technology raises exciting possibilities for detector doping at unprecedented levels with, for example, nanoparticles that include material with a high atomic number to increase both light scattering and the photoelectric absorption for gamma-ray detection [Taylor2011NanofluidCollectors, Mutreja2023HighAspects].
The third category of improvements relates to improving particle identification and topological reconstruction performance. Machine learning for event topological reconstruction could be appropriate for this application. One such method is to train a neural network by simulated data to generate maximum likelihood estimations of event characteristics given a series of observables [Eller2023APrinciples] (in this case, the number of photoelectrons in each channel). Lastly, the base solvent used in oWbLS (DIN) [Paweczak2013StudiesMethod] as well as transparent water-based liquid scintillators [Ford2022Pulse-shapeScintillators] have been shown to have the ability to discriminate between neutron and gamma-ray interactions based on the difference in the resulting pulse shape. The feasibility of this approach in the opaque regime should be investigated.